OpenStudy (anonymous):

write equation to solve using whole numbers and the distributive property. 8 bus rides that cost $2.65 each and 4 waters at $1.45 each. how much will it cost to ride and buy water.

OpenStudy (anonymous):

8(2.65) + 4(1.45) = 21.20 + 5.80 = 27.00